OpenStudy (anonymous):

Graph the Inequality 2x+5y<10

OpenStudy (mathstudent55):

First, look at the corresponding equation, 2x + 5y = 10 Can you graph that line?

OpenStudy (anonymous):

Yeah I think

OpenStudy (mathstudent55):

Ok, then graph the equation using a dashed line. You use a dashed line in an inequality that has < or >. If the inequality has <= or >=, then you use a solid line.

OpenStudy (anonymous):

can you guide me through the first step i forgot

OpenStudy (mathstudent55):

After you graph the corresponding equation, pick a point anywhere on the coordinate plane but not on the line. If (0, 0) is not on the line, pick (0, 0) since it's easy to work with. Then see if (0, 0) is part of the solution of the inequality by substituting it into the inequality. If it makes the inequality true, it is part of the solution. If it is part of the solution, then shade the side of the line that contains point (0,0). If not, then shade the other side.

OpenStudy (mathstudent55):

Now if you need more help, we can go through this step by step, but first read my previous answers where the entire graphing procedure is explained.

OpenStudy (anonymous):

yes i do need help, which i get about the shading but i forgot how to get the cordinates

OpenStudy (anonymous):

right

OpenStudy (mathstudent55):

Ok. Let's start from the beginning. To graph the inequality, 2x+5y<10, you must first graph the equation, 2x+5y=10, using a dashed line.

OpenStudy (mathstudent55):

The equation 2x+5y=10, represents a line. Let's graph it using the intercpets. We make a table to get two points. Since it's a line, two points are all you need.

OpenStudy (mathstudent55):